# Tiled Diffusion

### v1

@@ Abstract
For additional materials and source code, please visit our paper webpage
@@ 2 Related Work
We begin by discussing classical tiling techniques, highlighting the evolution of this field. We then explore recent advancements in image generation models, focusing on those applicable to image synthesis and editing. Finally, we examine current approaches to automated tiling using generative models, emphasizing the gaps our work aims to address.
@@ 2.3 Automated Tiling with Generative Models
Recent approaches have attempted to leverage generative models for automated tiling, but none have achieved the flexibility and range of tiling scenarios that our method supports. ControlMat [27] and TexTile [19] focus on self-tiling texture generation. (Textures are images with repetitive details or patterns with no high-level structure.) While these methods show promise in their specific domain, they cannot synthesize more general, non-texture, results. Additionally, they are limited to self-tiling scenarios.
Concurrent with our work, Content-aware Tile Generation [22] extends self-tiling to many-to-many tiling (Wang tiles) but remains constrained to texture synthesis and uses a multi-stage process with separate inpainting steps. Similarly, Seamless Tile Inpainting (STI) [3] uses inpainting diffusion models to create seamless connections between swapped image quarters for self-tiling, or adjacent images for one-to-one connections. These methods may suffer from inconsistencies because inpainting is a post processing-step, which cannot influence the initial generation of the images to make them fit together semantically. In contrast, our approach works directly on the latent representations during image generation. By creating all the different tiles simultaneously, we can share the necessary tiling information between the images being generated. This process makes our results inherently tileable across various domains, reducing inconsistencies and ensuring seamless connections. Asymmetric Tiling (AT) [26] modifies the Stable Diffusion architecture by replacing standard padding with circular padding in the Text encoder, UNet and VAE. This integration of tiling awareness into the diffusion process produces inherently tileable images. However, it is restricted to self-tiling scenarios with no rotations.
@@ 3.2.3 Constraint Application Process
After applying all the constraints throughout the diffusion process, we decode the latent representations and crop the results to the original image dimensions, producing seamlessly tiled images that satisfy all predefined constraints.
@@ 4.2 Quantitative Evaluation
We propose TS to measure tiling quality across various image domains. While existing methods for assessing tiling quality, like TexTile [19], focus on textures and are unable to generalize to other domains, our method is applicable to a broader range of images (examples in supplementary material). TS is calculated as the average of three separate measurements between two tiled images to reduce variance and capture potential discontinuities beyond the immediate connection area. Each measurement is defined as the mean absolute difference of pixel values at specific locations. Lower values indicate better tiling (range: 0.0 to 1.0). For a pair of images $I_{1}$ and $I_{2}$ tiled along an axis of length $h$:
@@
To measure image-text similarity (ITS), we use the cosine similarity between the CLIP encodings [17] of the image and the text. For perceptual similarity, we use LPIPS [30], which assesses the diversity of generated images compared to the VT2I images, with higher values indicating more diverse outputs. For overall image quality and realism, we use FID [8] between the generated images and VT2I. We also include a baseline evaluation for VT2I and Swapped VT2I, where we additionally calculate the same metrics for both. This baseline allows us to assess how our method performs relative to standard text-to-image generation. Table 1 shows that our method performs well across all metrics compared to other approaches. For this analysis, we evaluated tileability on the X-axis using the sampled prompts. For self-tiling, we created 1,000 constraints (one per prompt). For one-to-one tiling, we randomly formed 500 pairs of prompts, resulting in 500 constraints. In many-to-many scenarios, each constraint $C_{j}$ involved sets $A_{j}$ and $B_{j}$ with 2 sides each, using 4 prompts per constraint, totaling 250 constraints. Our Tiled Diffusion method performs consistently across all scenarios, including complex many-to-many tiling. The table also demonstrates the impact of removing key components of our method. Without the tiling constraint (TC), we achieve results similar to standard image generation — high quality but poor tiling. Without the similarity constraint (SC), performance remains stable except in many-to-many scenarios, where tiling quality decreases. This highlights the importance of both constraints in our method. Figure 5 illustrates how each metric (TS, ITS, LPIPS, FID) scales with increasing complexity in many-to-many tiling scenarios. Here, $n$ represents the number of sides within sets $A_{j}$ and $B_{j}$ for each constraint $C_{j}$, with $n=1$ corresponding to one-to-one tiling. As $n$ increases, the results remain nearly constant, demonstrating that our method scales effectively to larger sets of constraints while maintaining tileability and quality.
@@
Method Self-Tiling One-To-One Many-To-Many FID $\downarrow$ TS $\downarrow$ ITS $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ TS $\downarrow$ ITS $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ TS $\downarrow$ ITS $\uparrow$ LPIPS $\downarrow$ VT2I 0.29 0.31 0.29 0.31 0.29 0.31 Swapped VT2I 0.03 0.14 0.03 0.14 0.03 0.14 AT 49.2 0.03 0.29 0.79 STI 59.2 0.03 0.31 0.77 77.1 0.14 0.23 0.77 Tiled Diffusion 47.9 0.03 0.30 0.76 48.1 0.04 0.30 0.76 50.5 0.03 0.30 0.79 Tiled Diffusion w/o TC 48.1 0.28 0.30 0.77 48.4 0.31 0.30 0.77 50.9 0.30 0.30 0.79 Tiled Diffusion w/o SC 47.9 0.03 0.30 0.76 48.1 0.04 0.30 0.76 50.5 0.12 0.30 0.79
@@ 5.2 Tileable Texture Generation
Our Tiled Diffusion method creates wraparound panoramic images by generating seamlessly connected edges. This technique produces continuous views for wide-angle landscapes or architectural panoramas. The seamless edges ensure smooth transitions in wrapped formats.

### v4

@@ Abstract
For the supplemental material and code we refer the readers to our paper webpage
@@ 2 Related Work
We begin by discussing classical tiling techniques, highlighting the evolution of this field. We then explore recent advancements in image generation models, focusing on those applicable to image synthesis and editing. Next, we review tiling in popular applications. Finally, we examine current approaches to automated tiling using generative models, emphasizing the gaps our work aims to address.
@@ 2.2 Image Generation and Editing Models
2.3 Related Applications of Tiling
2.3.1 Texture Synthesis
There is vast research in creating tileable textures. Most recent approaches rely on input patches [24, 40, 39, 38, 22, 27, 23], multi-stage pipelines [40, 8, 27], or specialized training [38, 22, 37, 23]. Tiled Diffusion is unique in generating natural tileable textures directly from prompts using any latent diffusion model, not constrained to self-tiling scenarios.
2.3.2 Infinite/Panorama Image Generation
Unlike recent approaches for large-scale image generation [3, 35, 16], Tiled Diffusion generates a fixed set of tiles that can be seamlessly arranged in any order and quantity, offering a time- and space-efficient solution for infinite or panoramic imagery.
@@ 2.4 Automated Tiling with Generative Models
Seamless Tile Inpainting (STI) [4] uses inpainting diffusion models to create seamless connections between swapped image quarters for self-tiling, or adjacent images for one-to-one connections. Asymmetric Tiling (AT) [31] modifies the Stable Diffusion architecture by replacing standard padding with circular padding, producing inherently tileable images but restricted to self-tiling scenarios with no rotations.
In contrast, our approach works directly on the latent representations during image generation. By creating all the different tiles simultaneously, we can share the necessary tiling information between the images being generated. This process makes our results inherently tileable across various domains, reducing inconsistencies and ensuring seamless connections.
@@ 3.2.3 Constraint Application Process
After applying all the constraints throughout the diffusion process, we decode the latent representations and crop the results to the original image dimensions, producing seamlessly tiled images that satisfy all predefined constraints. We provide visual ablation of both constraints in the supplementary material.
@@ 4.2 Quantitative Evaluation
We propose TS to measure tiling quality across various image domains. While existing methods for assessing tiling quality, like TexTile [23], focus on textures and are unable to generalize to other domains, our method is applicable to a broader range of images (examples in supplementary material). TS is calculated as the maximum of three separate measurements between two tiled images to capture potential discontinuities beyond the immediate connection area. Each measurement is defined as the mean absolute difference of pixel values at specific locations. Lower values indicate better tiling (range: 0.0 to 1.0). For a pair of images $I_{1}$ and $I_{2}$ tiled along an axis of length $h$:
@@
To measure image-text similarity (CLIP-Score), we use the cosine similarity between the CLIP encodings [20] of the image and the text. For perceptual similarity, we use LPIPS [36], which assesses the diversity of generated images compared to the VT2I images, with higher values indicating more diverse outputs. For overall image quality and realism, we use FID [10] between the generated images and VT2I. We also include a baseline evaluation for VT2I and Swapped VT2I, where we additionally calculate the same metrics for both. This baseline allows us to assess how our method performs relative to standard text-to-image generation. Table 1 shows that our method performs well across all metrics compared to other approaches. For this analysis, we evaluated tileability on the X-axis using the sampled prompts. For self-tiling, we created 1,000 constraints (one per prompt). For one-to-one tiling, we randomly formed 500 pairs of prompts, resulting in 500 constraints. In many-to-many scenarios, each constraint $C_{j}$ involved sets $A_{j}$ and $B_{j}$ with 2 sides each, using 4 prompts per constraint, totaling 250 constraints. Our Tiled Diffusion method performs consistently across all scenarios. The table also demonstrates the impact of removing key components of our method. Without the tiling constraint (TC), we achieve results similar to standard image generation — high quality but poor tiling. Without the similarity constraint (SC), performance remains stable except in many-to-many scenarios, where tiling quality decreases. This highlights the importance of both constraints in our method. Figure 5 illustrates how each metric (TS, CLIP-Score, LPIPS, FID) scales with increasing complexity in many-to-many tiling scenarios. Here, $n$ represents the number of sides within sets $A_{j}$ and $B_{j}$ for each constraint $C_{j}$, with $n=1$ corresponding to one-to-one tiling. As $n$ increases, the results remain nearly constant, demonstrating that our method scales effectively to larger sets of constraints while maintaining tileability and quality.
Additionally, we provide texture synthesis analysis in the supplemental material where we compare our method to other methods in example-based texture synthesis, and evaluate our method in generating textures directly from prompts.
@@
Method Self-Tiling One-To-One Many-To-Many FID $\downarrow$ TS $\downarrow$ CLIP-Score $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ TS $\downarrow$ CLIP-Score $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ TS $\downarrow$ CLIP-Score $\uparrow$ LPIPS $\downarrow$ VT2I 0.29 0.31 0.29 0.31 0.29 0.31 Swapped VT2I 0.03 0.14 0.03 0.14 0.03 0.14 AT 49.2 0.03 0.29 0.79 STI 59.2 0.03 0.31 0.77 77.1 0.14 0.23 0.77 Tiled Diffusion 47.9 0.03 0.30 0.76 48.1 0.04 0.30 0.76 50.5 0.03 0.30 0.79 Tiled Diffusion w/o TC 48.1 0.28 0.30 0.77 48.4 0.31 0.30 0.77 50.9 0.30 0.30 0.79 Tiled Diffusion w/o SC 47.9 0.03 0.30 0.76 48.1 0.04 0.30 0.76 50.5 0.12 0.30 0.79
@@ 5.2 Tileable Texture Generation
Our Tiled Diffusion method creates wraparound panoramic images by generating seamlessly connected edges. This technique produces continuous views for wide-angle landscapes or architectural panoramas.